OpenStudy (anonymous):

Simplify 3 times f squared times g squared all over 12 times f to the eighth power.. pls help

OpenStudy (blurbendy):

( 3f^2 x g^2 ) / ( 12f^8) = g^2 / 4f^6 The f's on the top will cancel with two of the f's on the bottom which is how i got f^6 on the bottom g^2 stays the same on the top, and 3 / 12 = 4, which is how I got 4 on the bottom